The perimeter of square is 40m. The perimeter of rectangle of length 12m is as the same as the perimeter of square. Find which encloses the greater area square or rectangle?
Answer by nerdybill(7384)

The perimeter of square is 40m. The perimeter of rectangle of length 12m is as the same as the perimeter of square. Find which encloses the greater area square or rectangle?
Square:
since all four sides are equal:
40/4 = 10 m (length of one side)
area = 10*10 = 100 sq m
.
Rectangle:
Let w = width
then
2(w+12) = 40
w+12 = 20
w = 8 m
Area = 8*12 = 96 sq m
.
Answer: the square